OpenStudy (anonymous):

[Part 1 (4 points)] Show all work in dividing 4t^2-1/15t+10 / 4t^2-4t+1/3t^2+5t+2 [Part 2 (2 points)] Use complete sentences to explain your process.

OpenStudy (anonymous):

@Tsukiyama_Shuu

OpenStudy (tsukiyama_shuu):

Dude I suck at math...

OpenStudy (anonymous):

@jabez177

OpenStudy (lvana):

I'm a bit confused at to what it's supposed to look like. Can you take a screenshot?~

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

https://courses.vlacs.org/vlacontent/algebra2_v10_gs//assessments-xml/res0099/e_07_02_01aa.gif%22 follow that link and it has the equation on it

OpenStudy (anonymous):

did you see it?

OpenStudy (lvana):

Mhmm.~ It's quite the big problem, so I'll work it out for yah while explaining each step.~

OpenStudy (anonymous):

okay thank you. So far i have the you need to flip the fraction on the right side to get the recipricol and change the sign from division to multiplication

OpenStudy (lvana):

I managed to show you how to simplify part of this problem. I wanted to show you what I had gotten so far, so you wouldn't think I wasn't keeping you in the loop. Do tell me if you understand this.~

OpenStudy (anonymous):

Yes i understand it

OpenStudy (lvana):

Here is the next part.~

OpenStudy (anonymous):

@ivana you there?

OpenStudy (gabbyalicorn):

4t^2−1/15t+10/4t^2−4t+1/3t^2+5t+2 =4t^2+−115t+52t2+−4t+13t2+5t+2 Combine Like Terms =4t^2+−1/15t+5/2t2+−4t+1/3t2+5t+2 =(4t^2+52t2+13t2)+(−115t+−4t+5t)+(2) =41/6t2+14/15t+2 Answer: =41/6t^2+14/15t+2

OpenStudy (gabbyalicorn):

Please medal if you're satisfied with your answer. Thanks. :)

OpenStudy (anonymous):

can you walk me through each step? i dont understand what you did exactly @gabbyalicorn

OpenStudy (gabbyalicorn):

okay, can you wait a short while. Sorry. D:

OpenStudy (anonymous):

yes i can wait